OpenStudy (anonymous):

Let f(x)=(1+b^2)x^2+bx+1 and let m(b) be the minimum value of f(x).As b varies the range of m(b) is-

OpenStudy (anonymous):

ans. (0,1]

OpenStudy (anonymous):

remember the vertex formula for finding minimum of this quadratic function

OpenStudy (anonymous):

(-b/2a , -D/4a) ??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now u know that m(b)=-D/4a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well find the formula for m(b)

OpenStudy (anonymous):

could you pls solve!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

??

OpenStudy (anonymous):

sorry its not (0,1]

OpenStudy (anonymous):

@dpaInc help

OpenStudy (anonymous):

yes... i'm here.... :) @shubham.bagrecha, is (0, 1] the answer you got or is that the solution?

OpenStudy (anonymous):

thats solution

OpenStudy (anonymous):

hmmm. i keep getting (3/4, 1].... lemme try again...

OpenStudy (anonymous):

the vertex of the parabola is given by -b/(2a), so in this problem, the vertex is: \(\large \frac{-b}{2a}\rightarrow \frac{-b}{2(1+b^2)} \)

OpenStudy (anonymous):

U r Quite right answer is (3/4 ,1]

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

the y-coordinate is f(\(\large \frac{-b}{2(1+b^2)} \)) = \(\large \frac{3}{4}+\frac{1}{4(1+b^2)} \)

OpenStudy (anonymous):

as b varies, \(\large \frac{1}{4(1+b^2)} \) ranges from (0, 1/4]

OpenStudy (anonymous):

so the m(b) varies in the range (3/4,1]

OpenStudy (anonymous):

yes... but idk why the solution is (0, 1]...:(

OpenStudy (anonymous):

maybe question is f(x)=(1+b^2)x^2+2bx+1 .....

OpenStudy (anonymous):

@shubham.bagrecha if given function is f(x)=(1+b^2)x^2+bx+1 then answer is (3/4,1] not (0,1]